Two angles are complementary. One angle is 26 degrees more than the other angle. What ar the measures of the angles?
Answer by malakumar_kos@yahoo.com(315)


Two angles are complementary. One angle is 26 degrees more than the other angle. What ar the measures of the angles?
let one angle be = x its compliment is 90-x
given one angle is 26 degrees more than the other
therefore x = (90-x)+26
x+x = 90+26

2x = 116
x = 116/2 = 58
90-x = 90-58 = 32
the two angles are 58 degrees and 32 degrees